Develop 1/(1+z^2) in powers of z-a, a being a real number.Find the general coefficient and for a=1 reduce to simplest form.


1
Expert's answer
2021-11-10T14:25:45-0500

Taylor series:


"f(z)=\displaystyle{\sum_{k=0}^{\infin}}\frac{f^{(k)}(a)}{k!}(z-a)^k"


"f'(z)=-\frac{2z}{(1+z^2)^2}"


"f''(z)=\frac{6z^2-2}{(1+z^2)^3}"


"f'''(z)=\frac{24z(1-z^2)}{(1+z^2)^4}"


"\frac{1}{1+z^2}=\frac{1}{1+a^2}-\frac{2a}{(1+a^2)^2}(z-a)+\frac{6a^2-2}{2(1+a^2)^3}(z-a)^2+\frac{24a(1-a^2)}{6(1+a^2)^4}(z-a)^3+..."


for a=1:


"\frac{1}{1+z^2}=1-\frac{z}{2}+\frac{(z-1)^2}{4}-\frac{(z-1)^4}{8}+\frac{(z-1)^5}{8}-"